I have a problem that i have to solve that brings me to this: 5(26/25)+5y=-11.
I know that the following step is : 5y= -275/25-130/25= -405/25
y+-405/125 = - 81/25
What i want to know is where does -275 come from. I know that 5(26/25) is 130/25! but that is all. Please help!

Answer by Alan3354(69443)

5(26/25)+5y=-11.
I know that the following step is : 5y= -275/25-130/25= -405/25
y+-405/125 = - 81/25
What i want to know is where does -275 come from. I know that 5(26/25) is 130/25! but that is all. Please help!
----------------
-11 = -275/25
-------------------
I would do it like this:
5(26/25)+5y=-11
26/5 + 5y = -11
Eliminate fractions
26 + 25y = -55
25y = -81
y = -81/25